Figure 1: Some boxes and line styles Figure 2: Example schema level diagram

Figure 3: Example diagram of a tree structure Figure 4: Supertypes and subtypes Figure 5: A portion of a large model

Figure 6: Car model using EXPRESS-G Figure 7: Car model using Shlaer-Mellor

Figure 8: Car model using IDEF1X Figure 9: Car model using OMT Figure 10: Car model using E-R Figure 11: Car model using NIAM